OpenStudy (anonymous):

An electron rotates with the rate of 5*(10^14) cycle per second. Find intensity of the current produced due to that motion given that electron charge is 1.6*(10^-19) coulombs.

OpenStudy (anonymous):

do u knw the formula to calculate intensity of current ?

OpenStudy (anonymous):

I = Q/T or I = V/R

OpenStudy (anonymous):

yeah (Y)

OpenStudy (anonymous):

I don't have any of those, I only have the charge of the electron and how many times it rotates per second.

OpenStudy (anonymous):

that's the prob

OpenStudy (anonymous):

just multiply both the quantitites u will get the answer

OpenStudy (anonymous):

Yeah that's correct. But, is there a rule or any explanation as to why that is correct?

OpenStudy (anonymous):

actually u r applying ohms law !

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

If i find some explanation to this i will surely let u knw !